Please show me how to do these problems and please explains.
1.) A farmer has four sheepdogs and three beagles. If he randomly chooses a dog to accompany him on a walk, what is the probability of him taking a walk with sheepdog?
2.) Gordon spins a spinner with equal-sized sections numbered 1-6. In one spin, what is the likelihood that the spinner will stop on a 1 or a 5?
3.) When two number cubes labeled 1-6 are rolled, what is the probability that the result will be two 4's ?
4.) Joanne is guessing which day in November is Bess's birthday. Joanne know that Bess's birthday does not fall on an odd-numbered day. What is the probability that Joanne will guess the correct day on her first try?
Thank you for your help.

Answer by stanbon(75887)

Please show me how to do these problems and please explains.
1.) A farmer has four sheepdogs and three beagles. If he randomly chooses a dog to accompany him on a walk, what is the probability of him taking a walk with sheepdog?
Probability of event E = (# of ways E can happen)/(# of possible outcomes)
-----
P(sheepdog) = 4/7
---------------------------------

2.) Gordon spins a spinner with equal-sized sections numbered 1-6. In one spin, what is the likelihood that the spinner will stop on a 1 or a 5?
P(1 or 5) = 2/6 = 1/3
---------------------------------

3.) When two number cubes labeled 1-6 are rolled, what is the probability that the result will be two 4's ?
P(two 4's) = 1/36
--------------------------------
4.) Joanne is guessing which day in November is Bess's birthday. Joanne know that Bess's birthday does not fall on an odd-numbered day. What is the probability that Joanne will guess the correct day on her first try?
---
November has 30 days (15 even and 15 odd)
P(Joanne chooses the correct day) = 1/15
======================================================
Cheers,
Stan H.